-0.038 (18) T(18)= -4 +226
Answer #1

T(18) = - 4 + 22e-0.038(18)

= - 4 + 22e-0.684

= - 4 + 22*0.50459

= - 4 + 11.100

T(18) = 7.10108

Hope This Helps!